# Is the type-D NUT C-metric really “missing” from the most general Plebański-Demiański solution?

### v1

@@ II.3 ISM construction of the accelerating Kerr-NUT solution
The generated metric represents the accelerating Kerr-NUT solution, whose line element reads
@@
When $C_{1}=C_{2}=0$, the above generated 2-soliton solution just reduces to the C-metric (2).
Note that the above accelerating Kerr-NUT solution can also be regenerated via the “remove–re-add” tricky from the C-metric (2). The process is briefly sketched as follows. We first remove an anti-soliton $\nu_{1}$ at $z=z_{1}$ and a soliton $\mu_{2}$ at $z=z_{2}$ both with the trivial vectors $(1,0)$, then rescale the metric by a factor $\mu_{1}/\mu_{2}$. This essentially generates the above Rindler seed (20):
@@
Next we re-add the anti-soliton $\nu_{1}$ with the vector $m_{0}^{(1)}=(1,-C_{1}$) and the soliton $\mu_{2}$ with the vector $m_{0}^{(2)}=(1,C_{2}$), and multiple the generated metric by a factor $\mu_{2}/\mu_{1}$ so that the final metric is physics because it satisfies the canonical condition: $\det\mathbf{g}=-\rho^{2}$. The conformal factor can be easily evaluated via $f=f_{00}\det\mathbf{\Gamma}/{\det\mathbf{\Gamma}(C_{1}=0,C_{2}=0)}$, where $f_{00}$ is the conformal factor given by Eq. (2).
@@ II.4 Astorino’s parametrization
We first transform from Weyl-Lewis-Papapetrou coordinates ($\rho,z$) to spherical-like coordinates ($r,x$) via:
@@
It is very tedious to convert the four functions ($A,B,C,H$) into the expressions in terms of spherical-like coordinates. However, if we take
@@
then we are able to first write the 2-dimensional metric as
@@
After rescaling the Killing coordinates as $t\mapsto\sqrt{\alpha}t\,,\varphi\mapsto\varphi/\sqrt{\alpha}$, the above two parts are collectively written in terms of Lewis-Papapetrou coordinates as PRD109-084038
@@
which can also be recast into Boyer-Lindquist-like form
@@
When the accelerating parameter is set to zero, the above type-D NUT C-metric apparently reduces to the Taub-NUT solution, and also it obviously recovers the usual C-metric when the NUT charge vanishes. There is a natural question remained, namely, whether it is superficially accelerating just as pointed out in Ref. CQG22-3467 for the traditional form. Then, if it is so, we still lack a type-D NUT C-metric. However, this case seems unlikely. So, we will assume that the above NUT C-metric is not isometric to the Taub-NUT solution in the remaining context, but clearly an obvious proof is still needed.
The above solution has also been generalized in Ref. 2404.06551 to include the electric and magnetic charges as well as a nonzero cosmological constant. However, it suffices enough for our aim to turn off these three parameters. Also, in our opinion, it becomes useful only in the non-rotating case, as we will demonstrate the above rotating and accelerating solution can be transformed into the traditional form of the general P-D metric.
II.5 Griffiths-Podolsky’s parametrization
We next transform from Weyl-Lewis-Papapetrou coordinates ($\rho,z$) to spherical-like coordinates ($\tilde{r},\tilde{x}$) via:
@@ II.5 Griffiths-Podolsky’s parametrization
It is troublesome to convert the four functions ($A,B,C,H$) into the expressions in terms of spherical-like coordinates. A simple trick to do this relatively easy is to avoid using the square-root, such as, exploiting identities: $\tilde{\sigma}^{2k}=\big{(}\tilde{m}^{2}+\tilde{n}^{2}-\tilde{a}^{2}\big{)}^{k}$ and $\tilde{\sigma}^{2k+1}=\big{(}\tilde{m}^{2}+\tilde{n}^{2}-\tilde{a}^{2}\big{)}^{k}\tilde{\sigma}$ can efficiently facilitate to rationalize the expressions containing the square-root. This experience is also very effective to manipulate the phantom soliton whose square-root function could not be reduced to a simple algebraic expression.
@@
One can note that the above metric resembles somewhat like the one derived in the last subsection. In particular, they only differ by an overall constant factor in the case without the NUT charge and by further making identifications: $\alpha=\tilde{\alpha}\tilde{a}$, $m=\tilde{m}$, $a=\tilde{a}$, $r=\tilde{r}$, $x=\tilde{x}$.
@@
To summarize, we have re-derived the accelerating Kerr-NUT solution via the ISM procedure from the Rindler background seed. Then using two different parameterizations, we have recast the solution into two distinct forms in terms of spherical-like coordinates, namely, the one novelly given by Ref. PRD109-084038 , and another by the traditional one. Since the latter is already known to be of type-D, so is the one delivered by Astorino in PRD109-084038 . However, one still needs to find what coordinate transformations that may relate them. We leave this subject to the next section.
III Möbius transformations amongst two metrics
In order to find the relation between the above two metrics, a direct routine is to examine the solutions given by Eqs. (33) and (41), respectively, since both of them adopt the same Killing coordinates ($t,\varphi$). Therefore, the remaining most possible coordinate transformation is the Möbius transformations amongst the radial and azimuthal coordinates, respectively, together with further parameter identifications.
Let’s consider the following Möbius transformations:
@@ III Möbius transformations amongst two metrics
We first apply the above transformations to $\widetilde{\Sigma}=\tilde{r}^{2}+(\tilde{n}+\tilde{a}\tilde{x})^{2}$ and find its relation to $\Sigma$. To do so, it is easy to find that if the identity $\widetilde{\Sigma}(1-\nu\,x)^{2}[\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)]^{2}=C\,\Sigma$ holds true, then we must have the following conditions:
@@
Next, we seek the links between the 1-forms ($\tilde{X},\tilde{Y}$) in (42) and their counterparts ($X,Y$) without rescaling the Killing coordinates:
@@
Note that the relation for $\alpha^{2}$ listed above then becomes the identity for $\tilde{\alpha}^{2}$, so $\alpha$ can now temporarily be viewed as a free parameter too. Alternatively, we can also express this in term of untilded parameters as follows:
@@
Our next step is to find the relation between the mass parameters. Using the following identities:
@@
Finally, with the help of the following useful expressions:
@@
$\frac{\tilde{\alpha}^{2}}{\alpha^{2}}=\frac{\big{(}\nu^{4}+1\big{)}\tilde{a}^{2}-\nu\big{(}3\nu^{2}+1\big{)}\tilde{a}\tilde{n}+\big{(}3\nu^{2}-1\big{)}\tilde{n}^{2}}{\tilde{a}(\tilde{a}-\nu\tilde{n})\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}\big{(}1-\nu^{2}\big{)}}\,,$ (53)
@@
One can also conduct a similar task as done in the above by using the C-metric-like coordinates ($x,y$) with $y=-1/(\alpha\,r)$ and ($\bar{x},\bar{y}$) defined via $\bar{y}=-1/(\tilde{\alpha}\tilde{r})$, $\bar{x}=\tilde{n}+\tilde{a}\tilde{x}$, rather than by employing the spherical-like coordinates. In this case, both metrics become more symmetric, and now $\tilde{a}$ behaves like a scale parameter, for example, $\bar{y}=\tilde{n}+\tilde{a}(y-\nu)/(1-\nu\,y)$. Since the results remain the same as those given in the above, we will not repeat this work although it might be relatively easy.
@@
Previously, Griffiths and Podolsky CQG22-3467 demonstrated that in the nonrotating limit, the NUT C-metric is apparently accelerating and can be transformed into the Taub-NUT solution, as a result of this, the type-D NUT C-metric seems to be absent from the most general family of the type-D P-D solution, leaving a puzzle for almost two decades. The reason for this is probably that they didn’t consider the Möbius transformation. On the other hand, it is a simple matter to just set $a=0$ in the Astorino’s expression PRD109-084038 to arrive at the type-D C-metric, but this still does not explain how to get the type-D NUT C-metric from the most general P-D solution.
To educate the lesson why the type-D NUT C-metric is missing from the P-D metric, let’s consider what limit we should take in the original P-D metric so as to get the expected type-D NUT C-metric. Let’s first assume that the transformation parameter $\nu$ and all the solution parameters ($\tilde{m},\tilde{n},\tilde{a},\tilde{\alpha}$) as well as the radial coordinate $\tilde{r}$ are all finite and nonzero (although $\tilde{a}$ can be set to zero). To achieve the non-rotating NUT C-metric, we must first let the rotation parameter $a$ tend to zero ($a\to 0$). Then from the first identity in (45), the NUT charge parameter $n$ must also tend to a very tiny constant ($n\simeq 0$) so that $n/a$ keeps finite. From the fourth one in (45) or (47), we know that the acceleration parameter $\alpha$ must approach the infinity ($\alpha\to\infty$) so that $\alpha\,a$ is also finite. The second identity in (45) tells us that the overall factor $C$ should tend to the infinity ($C\to\infty$) so that $C\alpha$ can be kept finite. From the inverse Möbius transformation (44), we can recognize that the radial coordinate $r$ must approach zero so that $\alpha\,r$ remains finite. Finally, from the second Eq. (50), we learn that the mass parameter $m$ should tend to a very tiny constant $(m\simeq 0)$. In the $a\to 0$ limit, both ($\alpha\,a$, $\alpha\,n$, $\alpha\,m$) and $\alpha\,r$ remains fixed.
@@
apart from the appropriate rescaling of the whole line element plus the linear combinations of the Killing coordinates ($t\,,\varphi$). However, the key measure is to consider the above Möbius transformations (43).
@@ IV Transforming the type-D NUT C-metric into the familiar P-D form
In this section, we will consider another obvious coordinate transformations that bring the type-D NUT C-metric (35) into the familiar P-D form. To this end, let’s introduce the following coordinate transformations:
@@
which bring the NUT C-metric (35) into the P-D form:
@@
in which $\alpha$ can be simply set to unity. The only nonzero Kretschmann invariant is:
@@ V Concluding remarks
In this paper, we have shown that the type-D accelerating Kerr-NUT metric presented in Ref. PRD109-084038 is not a new solution, but it is still very useful in the non-rotating case, namely, the type-D static accelerating solution with a nonzero NUT charge since this solution is unavailable in the previous literature. We have justified this point from two aspects. First, we have applied the ISM procedure to generate the accelerating Kerr-NUT solution from the seed metric – Rindler vacuum background. Then we adopted two different parameterizations to derive the metric given by Astorino in Ref. PRD109-084038 and the traditional one early delivered by Griffiths and Podolsky. This clearly demonstrates that these two metrics are isometric, though displayed in two different forms. Second, we have also provided the obvious coordinate transformations (Möbius transformations of the radial and angular coordinates and linear combinations of the Killing coordinates) and parameter identifications between these two line elements. In particular, we also presented another concrete Möbius transformation and the linear combinations of the Killing coordinates that obviously cast the type-D NUT C-metric into the familiar form of the P-D solution. We anticipate that this work can give a final answer to resolve the “missing” puzzle of the type-D NUT C-metric.
@@
In the appendix Appendix: ISM construction of 5-dimensional static solutions with our new normalization scheme, we have given two simple examples to illustrate our new normalization scheme by deriving the 5-dimensional Schwarzschild-Tangherlini solution from the flat vacuum via the ISM construction. Our normalization strategy can be viewed as a methodological complementary to the Pomeransky’s “remove–re-add” tricky.
Note added: When this paper is under review (while arXiv on-holds our submission on August 27), an eprint 2409.02308 appears with a partial overlap of the subject for which the Möbius transformation is also used to relate different forms of the accelerating Kerr-NUT solution. The Möbius transformation previously used in Ref. PRD93-044021 was suggested to Dr. M. Astorino by S.-Q. Wu in an email communication about five months ago.
Acknowledgements.

### v3

@@ II.3 ISM construction of the accelerating Kerr-NUT solution
Then the generated metric represents the accelerating Kerr-NUT solution, whose line element reads
@@
When $C_{1}=C_{2}=0$, the above generated 2-soliton solution just reduces to the C-metric (2). So according to the perspective in Sec. II.1, the generated accelerating Kerr-NUT solution (26) can be interpreted as a Kerr-NUT black hole sitting on the Rindler background.
Note that the above accelerating Kerr-NUT solution (26) can also be regenerated via the “remove–re-add” tricky from the C-metric (2). The process is briefly sketched as follows. We first remove an antisoliton $\nu_{1}$ at $z=z_{1}$ and a soliton $\mu_{2}$ at $z=z_{2}$ both with the trivial vectors $(1,0)$, then rescale the metric by a factor $\mu_{1}/\mu_{2}$. This essentially generates the above Rindler seed (20):
@@
Next we re-add the antisoliton $\nu_{1}$ with the vector $m_{0}^{(1)}=(1,-C_{1}$) and the soliton $\mu_{2}$ with the vector $m_{0}^{(2)}=(1,C_{2}$), and multiply the generated metric by a factor $\mu_{2}/\mu_{1}$ so that the final metric is physics because it satisfies the canonical condition: $\det\mathbf{g}=-\rho^{2}$. The conformal factor can be easily evaluated via $f=f_{00}\det\mathbf{\Gamma}/{\det\mathbf{\Gamma}(C_{1}=0,C_{2}=0)}$, where $f_{00}$ is the conformal factor given by Eq. (2).
@@ II.4 Astorino’s parametrization
We first transform from the Weyl-Lewis-Papapetrou coordinates ($\rho,z$) to the spherical-like coordinates ($r,x$) via
@@
It is very tedious to convert the four functions ($A,B,C,H$) into the expressions in terms of the spherical-like coordinates. However, if we take
@@
then we are able to first write the two-dimensional metric as
@@
After rescaling the Killing coordinates as $t\mapsto\sqrt{\alpha}t\,,\varphi\mapsto\varphi/\sqrt{\alpha}$, the above two parts are collectively written in terms of the Lewis-Papapetrou coordinates as PRD109-084038
@@
which can also be recast into the Boyer-Lindquist-like form
@@
In Sec. III (especially Sec. IIIB) of Ref. PRD109-084038 , the above metric (32) was derived through an equivalent limiting procedure ($a_{2}\to 0,l_{2}\to 0$ and then $w_{4}\to\infty$) on the double Kerr-NUT solution. It was also claimed there that this spacetime is of type D, and might be included into the P-D family of black holes. Here we have built this metric with two solitons on a Rindler accelerating background. In the subsequent subsections, we will give an explicit verification of the Astorino’s insight by showing that the solution (33) is indeed diffeomorphic to the P-D metric.
@@
When the accelerating parameter is set to zero, the above type-D NUT C-metric apparently reduces to the Taub-NUT solution, and also it obviously recovers the usual C-metric when the NUT charge vanishes. There is a natural question remaining, namely, whether it is superficially accelerating just as pointed out in Ref. CQG22-3467 for the traditional form. Then, if it is so, we still lack a type-D NUT C-metric. However, this case seems unlikely. So, we will assume that the above NUT C-metric is not isometric to the Taub-NUT solution in the remaining context, but clearly an obvious proof is still needed.
The above solution has also been generalized in Ref. 2404.06551 to include the electric and magnetic charges as well as a nonzero cosmological constant. However, it suffices enough for our aim to turn off these three parameters. Also, in our opinion, it becomes useful only in the static case (where the rotation parameter vanishes), as we will demonstrate the above rotating and accelerating solution can be related to the traditional form of the general P-D metric.
II.5 Griffiths-Podolský’s parametrization
We next transform from the Weyl-Lewis-Papapetrou coordinates ($\rho,z$) to the spherical-like coordinates ($\tilde{r},\tilde{x}$) via
@@ II.5 Griffiths-Podolský’s parametrization
It is troublesome to convert the four functions ($A,B,C,H$) into the expressions in terms of the spherical-like coordinates. A simple trick to do this relatively easily is to avoid using the square root, such as, exploiting identities: $\tilde{\sigma}^{2k}=\big{(}\tilde{m}^{2}+\tilde{n}^{2}-\tilde{a}^{2}\big{)}^{k}$ and $\tilde{\sigma}^{2k+1}=\big{(}\tilde{m}^{2}+\tilde{n}^{2}-\tilde{a}^{2}\big{)}^{k}\tilde{\sigma}$ can efficiently facilitate to rationalize the expressions containing the square root. This experience is also very effective to manipulate the phantom soliton whose square-root function could not be reduced to a simple algebraic expression.
@@
One can note that the above metric resembles somewhat the one derived in the last subsection. In particular, they only differ by an overall constant factor in the case without the NUT charge and by further making identifications: $\alpha=\tilde{\alpha}\tilde{a}$, $m=\tilde{m}$, $a=\tilde{a}$, $r=\tilde{r}$, $x=\tilde{x}$.
@@
To summarize, we have rederived the accelerating Kerr-NUT solution via the ISM procedure from the Rindler background seed. Then using two different parametrizations, we have recast the solution into two distinct forms in terms of the spherical-like coordinates, namely, the one novelly given by Ref. PRD109-084038 , and another by the traditional one. Since the latter is already known to be of type D, so is the one delivered by Astorino in PRD109-084038 . However, one still needs to find what coordinate transformations may relate them. We leave this subject to the next section.
By the way, here we would like to mention the related work PRD93-044021 that dealt with the Euclidean P-D metric. In the Sec. VC of that paper, Chen obtained the Euclidean accelerating Kerr-NUT metric as a 3-soliton solution via the ISM procedure. Then he performed a Möbius transformation on the coordinates ($u,v$) and redefined the Killing coordinates, and further made various parameter identifications so that the 3-soliton solution is brought to the Euclidean P-D solution in the familiar form. This hints that the Möbius transformation may be relevant to the present work.
III Möbius transformation amongst two metrics
In order to find the relation between the above two metrics, a direct routine is to examine the solutions given by Eqs. (33) and (41), respectively, since both of them adopt the same Killing coordinates ($t,\varphi$). Therefore, the remaining most possible coordinate transformation is the Möbius transformation amongst the radial and azimuthal coordinates, respectively, together with further parameter identifications.
Let us consider the following Möbius transformation:
@@ III Möbius transformation amongst two metrics
We first apply the above transformation to $\widetilde{\Sigma}=\tilde{r}^{2}+(\tilde{n}+\tilde{a}\tilde{x})^{2}$ and find its relation to $\Sigma$. To do so, it is easy to find that if the identity $\widetilde{\Sigma}(1-\nu\,x)^{2}[\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)]^{2}=C\,\Sigma$ holds true, then we must have the following conditions:
@@
Next, we seek the links between the one-forms ($\tilde{X},\tilde{Y}$) in (42) and their counterparts ($X,Y$) without rescaling the Killing coordinates:
@@
Note that the relation for $\alpha^{2}$ listed above then becomes the identity for $\tilde{\alpha}^{2}$, so $\alpha$ can now temporarily be viewed as a free parameter too. Alternatively, we can also express this in terms of untilded parameters as follows:
@@
Our next step is to find the relation between the mass parameters. Using the following identities,
@@
Finally, with the help of the following useful expressions,
@@
$\frac{\tilde{\alpha}^{2}}{\alpha^{2}}=\frac{\big{(}\nu^{4}+1\big{)}\tilde{a}^{2}-\nu\big{(}3\nu^{2}+1\big{)}\tilde{a}\tilde{n}+\big{(}3\nu^{2}-1\big{)}\tilde{n}^{2}}{\tilde{a}(\tilde{a}-\nu\tilde{n})\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}\big{(}1-\nu^{2}\big{)}}\,.$ (53)
@@
One can also conduct a similar task as done in the above by using the C-metric-like coordinates ($x,y$) with $y=-1/(\alpha\,r)$ and ($\bar{x},\bar{y}$) defined via $\bar{y}=-1/(\tilde{\alpha}\tilde{r})$, $\bar{x}=\tilde{n}+\tilde{a}\tilde{x}$, rather than by employing the spherical-like coordinates. In this case, both metrics become more symmetric, and now $\tilde{a}$ behaves like a scale parameter; for example, $\bar{y}=\tilde{n}+\tilde{a}(y-\nu)/(1-\nu\,y)$. Since the results remain the same as those given in the above, we will not repeat this work although it might be relatively easy.
@@
Previously, Griffiths and Podolský CQG22-3467 demonstrated that in the limit of zero angular momentum parameter, the NUT C-metric is apparently accelerating and can be transformed into the Taub-NUT solution; as a result of this, the type-D NUT C-metric seems to be absent from the most general family of the type-D P-D solution, leaving a puzzle for almost two decades. The reason for this is probably that they did not consider the Möbius transformation. On the other hand, it is a simple matter to just set $a=0$ in the Astorino’s expression PRD109-084038 to arrive at the type-D C-metric, but this still does not explain how to get the type-D NUT C-metric from the most general P-D solution.
To elucidate the lesson why the type-D NUT C-metric is missing from the P-D metric, let us consider what limit we should take in the original P-D metric to get the expected type-D NUT C-metric. Let us first assume that the transformation parameter $\nu$ and all the solution parameters ($\tilde{m},\tilde{n},\tilde{a},\tilde{\alpha}$) as well as the radial coordinate $\tilde{r}$ are all finite and nonzero (although $\tilde{a}$ can be set to zero). To achieve the static NUT C-metric, we must first let the rotation parameter $a$ tend to zero ($a\to 0$). Then from the first identity in (45), the NUT charge parameter $n$ must also tend to a very tiny constant ($n\simeq 0$) so that the ratio $n/a$ keeps finite. From the fourth one in (45) or (47), we know that the acceleration parameter $\alpha$ must approach the infinity ($\alpha\to\infty$) so that $\alpha\,a$ is also finite. The second identity in (45) tells us that the overall factor $C$ should tend to the infinity ($C\to\infty$) so that $C\alpha$ can be kept finite. From the inverse Möbius transformation (44), we can recognize that the radial coordinate $r$ must approach zero so that $\alpha\,r$ remains finite. Finally, from the second equation of Eq. (50), we learn that the mass parameter $m$ should tend to a very tiny constant $(m\simeq 0)$. In the $a\to 0$ limit, both ($\alpha\,a$, $\alpha\,n$, $\alpha\,m$) and $\alpha\,r$ remains fixed.
@@
apart from the appropriate rescaling of the whole line element plus the linear combinations of the Killing coordinates ($t\,,\varphi$). However, the key measure is to consider the above Möbius transformation (43).
@@ IV Transforming the type-D NUT C-metric into the familiar P-D form
In this section, we will consider another obvious coordinate transformation that brings the type-D NUT C-metric (35) into the familiar P-D form. To this end, let us introduce the following coordinate transformations:
@@
which bring the NUT C-metric (35) into the familiar P-D form:
@@
in which $\alpha$ can be simply set to unity. The only nonzero Kretschmann invariant is
@@ V Concluding remarks
In this paper, we have shown that the type-D accelerating Kerr-NUT metric presented in Ref. PRD109-084038 is not a new solution in the sense that it is diffeomorphic to the original P-D solution, but it is still very convenient for further investigations in the nonrotating case where the angular momentum parameter vanishes, namely, the type-D static accelerating solution with a nonzero NUT charge since this solution is unavailable in the previous literature. We have justified this point from two aspects. First, we have applied the ISM procedure to generate the accelerating Kerr-NUT solution from the seed metric – Rindler vacuum background, which can be equivalently obtained as in Ref. PRD109-084038 via a limiting procedure ($a_{2}\to 0,l_{2}\to 0$ and then $w_{4}\to\infty$) on the double Kerr-NUT solution. Then we adopted two different parametrizations to derive the novel metric given by Astorino in Ref. PRD109-084038 and the traditional one early delivered by Griffiths and Podolský. This clearly demonstrates that these two metrics are isometric, though displayed in two different forms. Second, we have also provided the obvious coordinate transformations (i.e., Möbius transformation of the radial and angular coordinates and linear combinations of the Killing coordinates) and parameter identifications between these two line elements. Therefore, our work provides an explicit verification of the insight of Ref. PRD109-084038 that the obtained spacetimes not only are of type D, but also might be included into the P-D family. In particular, we also presented another concrete Möbius transformation and the linear combinations of the Killing coordinates that obviously cast the type-D NUT C-metric into the familiar form of the P-D solution. We anticipate that this work can give a final answer to resolve the “missing” puzzle of the type-D NUT C-metric.
@@
In the Appendix Appendix: ISM construction of five-dimensional static solutions with our new normalization scheme, we have given two simple examples to illustrate our new normalization scheme by deriving the five-dimensional Schwarzschild-Tangherlini solution from the flat vacuum via the ISM construction. Our normalization strategy can be viewed as a methodological complement to the Pomeransky “remove–re-add” tricky.
Note added. Recently, an eprint 2409.02308 appeared with a partial overlap of the subject for which the Möbius transformation is also used to relate the Astorino’s new form to three different forms of the most general P-D solution. Our paper, however, focuses on the proof that the Astorino’s novel form can be reproduced from the already-known accelerating Kerr-NUT solution and explains why the type-D NUT C-metric is “missing” from the rotating and accelerating P-D solution.
Acknowledgements